OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (anonymous):

So, you want to put -11x + 9y = -12 in Slope-Intercept form?

OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (anonymous):

Do you know what Slope-Intercept form looks like?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

my daughter was helping me but she had to work

OpenStudy (anonymous):

Slope-Intercept form is\[\large y = mx + b \]Where "m" is the slope, and "b" is the y-intercept.

OpenStudy (anonymous):

Right now you're equation is in the Standard form, Ax + By = C. To change it into Slope-Intercept form, we need to solve for "y".

OpenStudy (anonymous):

\[ -11x + 9y = -12 \]Add 11x to both sides, and you have: \[9y = 11x - 12\]Now divide both sides by 9, to get the y alone.\[\frac{ 9y }{ 9 } = \frac{ 11 }{ 9 }x - \frac{ 12 }{ 9 }\]Simplify and you get:\[\Large y = 11/9x - 4/3\]

OpenStudy (anonymous):

\[ y = mx + b \]\[ y = 11/9x - 4/3 \] m(slope) = 11/9 b(y-intercept) = -4/3